# A flexible class of latent variable models for the analysis of antibody response data

### v1

@@ 6 Discussion
In this paper, we introduced a latent variable modelling framework that generalises the commonly used mixture model approach for the analysis of serological data. Classical Gaussian mixture models arise as a special case of our framework when the latent variable $T$ is restricted to a Bernoulli distribution. By modeling immune status as a continuous latent state rather than imposing discrete serostatus categories, the framework captures a broader range of immunological variation and avoids the artificial dichotomization inherent in standard finite mixture models. As illustrated in our application to malaria serology, this leads to a more flexible class of distributions that can accommodate the skewness and irregular shapes commonly observed in antibody data.
The primary contributions of this framework are threefold. First, it provides a principled way to model the full distribution of antibody concentrations without dichotomization, which remains the dominant practice in serological analyses (corran2007; Drakeley2005; Arnold2014; cox2022) despite well-established information loss (royston2006; kyomuhangi2021). By modelling antibody levels directly, the framework provides a more refined characterisation of how immune responses vary and evolve with age. Second, it offers a flexible approach for incorporating age-dependent structure through either mechanistic specifications that encode epidemiological assumptions about exposure and immune response dynamics, or data-driven specifications that adaptively capture empirical patterns without imposing strong structural constraints. Third, by explicitly representing immune activation as a continuum, the framework naturally accommodates complex population heterogeneity, such as distinguishing settings with bimodal seroreactivity (high proportions at both extremes) from those with intermediate or transitional immune states.
A further distinction must be made between mechanistic models applied to the latent variable $T$ in our framework and the same mechanistic models traditionally applied to dichotomised antibody outcomes. For example, when the reversible catalytic model in (11) is used for binary serostatus data, the parameter $\lambda$ is interpreted as the seroconversion rate, that is, the rate at which individuals transition from seronegative to seropositive status. In our framework, however, seroconversion is not a binary event, and $T$ represents a continuous underlying immune state. Consequently, the parameter $\lambda$ should be interpreted as the rate at which individuals develop stronger seroreactivity, which may or may not lead to a seropositive test outcome. This thus suggests that estimates of $\lambda$ obtained under our model are expected to be higher than those derived from dichotomised serological analyses. Another important feature of the framework is that it enables a more rigorous statistical validation of the biological assumptions underlying mechanistic models, an aspect that is rarely examined in standard serological analyses. For example, the application in Section 5 shows that a reversible catalytic model exhibits some, albeit slight, discrepancy with the data in the immune response dynamics in children below 6 years of age.
@@
Future research will also focus on extending the current modelling framework to address a wider range of questions in sero epidemiological research, including the development of joint models for multiple antibody responses. One natural direction is to replace the single latent quantity with a multivariate latent structure and to combine this with an appropriate multivariate Beta distribution, for example using the construction proposed by arnold2011. This would allow the model to capture dependencies between antibody responses to different antigens and provide deeper insight into how immune markers coevolve following exposure. The same idea extends naturally beyond serology, since the latent variable formulation is directly relevant to high dimensional omics settings such as proteomics and transcriptomics, where disease states are typically treated as discrete categories, for example healthy versus diseased or sepsis molecular endotypes. Recent work using neural network models for disease classification from multi omics data (hartman2023) and large transcriptomic and proteomic studies of sepsis subtypes (wong2012; reyes2020) all adopt this discrete label perspective, despite the underlying host response being biologically continuous rather than strictly partitioned. Within our framework this continuum can be represented by a latent variable $T$, with each biomarker, here each transcript or protein, modelled as an observed outcome $Y_{j}$ whose conditional distribution depends smoothly on $T$ through marker specific mean and variance functions that describe how expression evolves across the response axis. While a single latent variable provides a parsimonious representation of the dominant gradient of biological variation, many omics settings involve several distinct processes, for example inflammation, immune suppression, and endothelial injury, which may jointly determine the multivariate profile. In such cases it may be necessary to adopt a multivariate latent representation and to define the conditional distribution of $Y_{j}$ as a function of this collection of latent quantities, allowing the framework to capture more complex patterns and multiple dimensions of biological variation that cannot be reduced to a single continuum.
@@
In summary, the flexibility of the proposed latent variable modelling framework, together with its strong potential for further extensions, illustrate its broad applicability in sero epidemiological research. We believe that these strengths position it as a compelling approach for routine use in settings where a deeper understanding of immunological processes is of primary importance.
@@ Acknowledgements
We would like to thank Prof. Peter J. Diggle (Lancaster University), Dr. Irene Kyomuhangi (Lancaster University), Erik Hartman (Lund University), Prof. Joacim Rocklöv (University of Heidelberg), and Dr. Gillian Stresman (University of South Florida) for their valuable feedback, which helped to improve the manuscript.

### v2

@@ 7 Discussion
We have introduced a latent variable modelling framework that offers four key methodological contributions. First, it provides a principled way to model the full distribution of antibody concentrations without dichotomization, which remains the dominant practice in serological analyses (corran2007; Drakeley2005; Arnold2014; cox2022) despite well-established information loss (royston2006; kyomuhangi2021). By modelling antibody levels directly, the framework provides a more refined characterisation of how immune responses vary and evolve with age. Second, it offers a flexible approach for incorporating age-dependent structure through either mechanistic specifications that encode epidemiological assumptions about exposure and immune response dynamics, or data-driven specifications that adaptively capture empirical patterns without imposing strong structural constraints. Third, by explicitly representing immune activation as a continuum, the framework more naturally accommodates complex population heterogeneity in immunological responses. Fourth, we introduce an $L_{2}$ histogram-based estimator that provides a computationally efficient alternative to maximum likelihood estimation, enabling practical implementation of the framework even in computationally demanding contexts with large sample sizes.
A further distinction must be made between mechanistic models applied to the latent variable $T$ in our framework and the same mechanistic models traditionally applied to dichotomised antibody outcomes. For example, when the reversible catalytic model in (11) is used for binary serostatus data, the parameter $\lambda$ is interpreted as the seroconversion rate, that is, the rate at which individuals transition from seronegative to seropositive status. In our framework, however, seroconversion is not a binary event, and $T$ represents a continuous underlying immune state. Consequently, the parameter $\lambda$ should be interpreted as the rate at which individuals develop stronger seroreactivity, which may or may not lead to a seropositive test outcome. This thus suggests that estimates of $\lambda$ obtained under our model are expected to be higher than those derived from dichotomised serological analyses. Another important feature of the framework is that it enables a more rigorous statistical validation of the biological assumptions underlying mechanistic models, an aspect that is rarely examined in standard serological analyses. For instance, the application in Section 6 reveals subtle deviations from the reversible catalytic model in the immune response dynamics of children aged below 10 years. Such discrepancies would remain undetected in analyses based on dichotomized serostatus data.
@@
Our simulation study demonstrates that the $L_{2}$ estimator achieves speedup factors ranging from 7-fold at $n=100$ to over 250-fold at $n=5000$ relative to MLE, while maintaining reasonable statistical accuracy in most scenarios. However, MLE demonstrates superior asymptotic efficiency, particularly in challenging scenarios with symmetric or highly skewed latent distributions, where $L_{2}$ exhibits persistent bias even at large sample sizes. We therefore recommend a hybrid strategy for parameter estimation which consists of using the $L_{2}$ estimation for exploratory analysis and, whenever computationally feasible, refine final estimates using MLE with $L_{2}$ solutions as starting values. However in scenarios where there is a clear bimodal separation in the distribution of antibody levels, $L_{2}$ performs nearly as well as MLE even at moderate sample sizes and might be used without the need for further refinement of the estimates with MLE.
Future research will focus on extending the current modelling framework to address a wider range of questions in sero epidemiological research, including the development of joint models for multiple antibody responses. One natural direction is to replace the single latent quantity with a multivariate latent structure and to combine this with an appropriate multivariate Beta distribution, for example using the construction proposed by arnold2011. This would allow the model to capture dependencies between antibody responses to different antigens and provide deeper insight into how immune markers coevolve following exposure. The same idea extends naturally beyond serology, since the latent variable formulation is directly relevant to high dimensional omics settings such as proteomics and transcriptomics, where disease states are typically treated as discrete categories, for example healthy versus diseased or sepsis molecular endotypes. Recent work using neural network models for disease classification from multi omics data (hartman2023) and large transcriptomic and proteomic studies of sepsis subtypes (wong2012; reyes2020) all adopt this discrete label perspective, despite the underlying host response being biologically continuous rather than strictly partitioned. Within our framework this continuum can be represented by a latent variable $T$, with each biomarker, here each transcript or protein, modelled as an observed outcome $Y_{j}$ whose conditional distribution depends smoothly on $T$ through marker specific mean and variance functions that describe how expression evolves across the response axis. While a single latent variable provides a parsimonious representation of the dominant gradient of biological variation, many omics settings involve several distinct processes, for example inflammation, immune suppression, and endothelial injury, which may jointly determine the multivariate profile. In such cases it may be necessary to adopt a multivariate latent representation and to define the conditional distribution of $Y_{j}$ as a function of this collection of latent quantities, allowing the framework to capture more complex patterns and multiple dimensions of biological variation that cannot be reduced to a single continuum.
@@
In summary, the flexibility of the proposed latent variable modelling framework, together with its strong potential for further extensions, illustrate its broad applicability in sero epidemiological research and beyond. We believe that these strengths position it as a compelling approach for routine use in settings where a deeper understanding of immunological processes is of primary importance.
@@ Acknowledgements
We thank all those who contributed to the collection of data included in this paper, specifically the survey participants in Kenya, and the KEMRI/CDC research team.
@@
We would like to thank Prof. Peter J. Diggle (Lancaster University), Dr. Irene Kyomuhangi (Lancaster University), Ivan Hejný (Lund University), Erik Hartman (Lund University), Prof. Joacim Rocklöv (University of Heidelberg), and Dr. Gillian Stresman (University of South Florida) for their valuable feedback, which helped to improve the manuscript.
The computations described in this paper were performed using the University of Birmingham’s BlueBEAR HPC service, which provides a High Performance Computing service to the University’s research community. See www.birmingham.ac.uk/bear for more details.
Appendix A Proof of Theorem 4.1
Proof.
Let $f_{\vartheta,n}:=\;f_{n}(\cdot;\vartheta)$. We apply Corollary 3.2.3 of vandervaartwellner1996. This corollary establishes consistency provided that: (i) the empirical criterion converges uniformly to the population criterion, i.e., $\sup_{\vartheta\in\vartheta}|\mathbb{M}_{n}(\vartheta)-\mathbb{M}(\vartheta)|\xrightarrow{p}0$. and that (ii) the population maximum is well-separated. Assumption A2. directly gives condition (ii). Thus, it remains only to show condition (i).
We prove this condition in two step. We first establish that the unparameterized histogram $\widehat{f}_{n}$ converges to the true density $f_{0}$ in the $L_{2}$ norm. freedman1981histogram provide the exact mean integrated squared error (MISE) decomposition in their Proposition (1.10):
$\mathbb{E}\Bigl[\|\widehat{f}_{n}-f_{0}\|_{2}^{2}\Bigr]\;=\;\left(\frac{1}{nh_{n}}-\frac{1}{n}\|f_{0,n}\|_{2}^{2}\right)\;+\;\|f_{0,n}-f_{0}\|_{2}^{2},$
where $f_{0,n}$ is the theoretical discretized density (the expectation of $\widehat{f}_{n}$). The first term is bounded by $(nh_{n})^{-1}$ and under Assumption A4., this term vanishes. The second term corresponds to the approximation error of the binned density; as shown in Eq. (2.5) of freedman1981histogram, this vanishes as $h_{n}\to 0$ for any $f_{0}\in L_{2}$. Therefore, $\mathbb{E}[\|\widehat{f}_{n}-f_{0}\|_{2}^{2}]\to 0$, and by Markov’s inequality:
$\|\widehat{f}_{n}-f_{0}\|_{2}\;\xrightarrow{p}\;0.$ (16)
Using the first step we establish Uniform convergence of the criterion. Note that $|\mathbb{M}_{n}(\vartheta)-\mathbb{M}(\vartheta)|=|\|\widehat{f}_{n}-f_{\vartheta,n}\|_{2}^{2}-\|f_{0}-f(\cdot;\vartheta)\|_{2}^{2}|$. We decompose the difference as:
$|\mathbb{M}_{n}(\vartheta)-\mathbb{M}(\vartheta)|\;$ $\leq\;\Bigl|\|\widehat{f}_{n}-f_{\vartheta,n}\|_{2}^{2}-\|f_{0}-f_{\vartheta,n}\|_{2}^{2}\Bigr|$
$\quad+\;\Bigl|\|f_{0}-f_{\vartheta,n}\|_{2}^{2}-\|f_{0}-f(\cdot;\vartheta)\|_{2}^{2}\Bigr|.$
We prove separately that the two terms on the right goes to zero. By the triangle inequality for norms, $|\|a\|_{2}^{2}-\|b\|_{2}^{2}|\leq\|a-b\|_{2}(\|a\|_{2}+\|b\|_{2})$. Applying the inequality to the first term (with $a=\widehat{f}_{n}-f_{\vartheta,n}$ and $b=f_{0}-f_{\vartheta,n}$):
$\Bigl|\|\widehat{f}_{n}-f_{\vartheta,n}\|_{2}^{2}-\|f_{0}-f_{\vartheta,n}\|_{2}^{2}\Bigr|\;\leq\;\|\widehat{f}_{n}-f_{0}\|_{2}\left(\|\widehat{f}_{n}-f_{\vartheta,n}\|_{2}+\|f_{0}-f_{\vartheta,n}\|_{2}\right).$
By Assumption A3., the terms in the parenthesis are bounded uniformly in probability. Thus by (16), $\Bigl|\|\widehat{f}_{n}-f_{\vartheta,n}\|_{2}^{2}-\|f_{0}-f_{\vartheta,n}\|_{2}^{2}\Bigr|$ converges to 0 uniformly in $\vartheta$.
Applying the inequality to the second term:
$\Bigl|\|f_{0}-f_{\vartheta,n}\|_{2}^{2}-\|f_{0}-f(\cdot;\vartheta)\|_{2}^{2}\Bigr|\;\leq\;\|f_{\vartheta,n}-f(\cdot;\vartheta)\|_{2}\left(\|f_{0}-f_{\vartheta,n}\|_{2}+\|f_{0}-f(\cdot;\vartheta)\|_{2}\right).$
Again, following Assumption A3. and (16), the terms in parentheses are bounded uniformly and bounded in probability (note that if two terms tend to zero in probability, so does their product). freedman1981histogram show in Proposition (2.7) that when the density is absolutely continuous with square-integrable derivative, then $\|f_{\vartheta,n}-f(\cdot;\vartheta)\|^{2}_{2}$ is $O(h_{n}^{2})$. Specifically, for small $h_{n}$, the term is bounded by a constant times $h_{n}^{2}\|f^{\prime}(\cdot;\vartheta)\|_{2}^{2}$. Thus under Assumption A5., $\sup_{\vartheta\in\vartheta}\|f^{\prime}(\cdot;\vartheta)\|_{2}<\infty$:
$\sup_{\vartheta\in\vartheta}\|f_{\vartheta,n}-f(\cdot;\vartheta)\|_{2}\;\leq\;Ch_{n}\sup_{\vartheta\in\vartheta}\|f^{\prime}(\cdot;\vartheta)\|_{2}\;\longrightarrow\;0.$
Thus, $\sup_{\vartheta\in\vartheta}|\mathbb{M}_{n}(\vartheta)-\mathbb{M}(\vartheta)|\xrightarrow{p}0$.
Appendix B Verification of regularity conditions for the latent variable model
To apply the consistency result of Theorem 4.1 to our proposed latent variable model, we must verify that the model defined in (2) satisfies Assumptions A1A5.
Lemma B.1.
Consider the latent variable model
$f(y;\vartheta)\;=\;\int_{0}^{1}\phi(y;\mu(t),\sigma^{2}(t))\,g_{T}(t;\psi)\,dt,$
where $\mu(t)=(1-t)\mu_{0}+t\mu_{1}$ and $\sigma^{2}(t)=(1-t)\sigma_{0}^{2}+t\sigma_{1}^{2}$. Assume the parameter space $\vartheta$ is compact (A1) and satisfies the ordering constraint $\mu_{0}<\mu_{1}$ to avoid label switching. Furthermore, assume $\sigma_{0},\sigma_{1}\geq c$ for some constant $c>0$ to ensure the density is bounded. Let $g_{T}$ be identifiable and continous with respect $\psi$. a If the histogram discretization satisfies A4, then the estimator satisfies Assumptions A1A5.
Proof.
We verify each assumption separately:
  1. A1.
    Compactness. Satisfied by the definition of the parameter space $\vartheta$.
  2. A2.
    Identification. The condition requiring a well-separated maximum is satisfied by the combination of model identifiability, continuity, and compactness.
    First, the mapping $\vartheta\mapsto f(\cdot;\vartheta)$ is injective, ensuring that the population criterion $\mathbb{M}(\vartheta)$ has a unique maximizer $\vartheta_{0}$. This injectivity holds because the $g_{T}(t;\psi)$ is identifiable, and the ordering constraint $\mu_{0}<\mu_{1}$ ensures a strictly monotonic (bijective) mapping from the latent variable to the conditional moments.
    Second, the criterion $\mathbb{M}(\vartheta)$ is continuous on $\vartheta$ because both the Gaussian kernel $\phi$ and the mixing density $g_{T}$ are continuous functions of the parameters. To see that the maximum is well-separated, consider any open neighborhood $G$ of $\vartheta_{0}$. The set $K=\vartheta\setminus G$ is a closed subset of a compact space, and thus is itself compact. By the Extreme Value Theorem, the continuous function $\mathbb{M}$ must attain a maximum on $K$. Since $\vartheta_{0}$ is the unique maximizer and $\vartheta_{0}\notin K$, this attained maximum must be strictly less than $\mathbb{M}(\vartheta_{0})$:
    $\sup_{\vartheta\in\vartheta\setminus G}\mathbb{M}(\vartheta)\;=\;\max_{\vartheta\in K}\mathbb{M}(\vartheta)\;<\;\mathbb{M}(\vartheta_{0}).$
  3. A3.
    Boundedness. We must show $f(y;\vartheta)$ is uniformly bounded. The Gaussian kernel $\phi(y;\mu,\sigma^{2})$ is bounded by $(2\pi\sigma^{2})^{-1/2}$. By the assumption that component variances are bounded away from zero ($\sigma_{0},\sigma_{1}\geq c$), the interpolated variance satisfies $\sigma(t)\geq c$ for all $t\in[0,1]$. Thus,
    $f(y;\vartheta)\;=\;\int_{0}^{1}\phi(y;\mu(t),\sigma^{2}(t))g_{T}(t;\psi)\,dt\;\leq\;\frac{1}{\sqrt{2\pi}c}\int_{0}^{1}g_{T}(t;\psi)\,dt\;=\;\frac{1}{\sqrt{2\pi}c}.$
    Since $f$ is bounded by a constant $K$ and $\int f(y)dy=1$, it follows that $\|f\|_{2}^{2}\leq K\int f(y)dy=K<\infty$, so $f\in L_{2}$.
  4. A4.
    Histogram Regime. This property can be easily satisfied by choosing any histogram binning rule for which $h_{n}\to 0$ and $nh_{n}\to\infty$, such as the Sturges rule adopted in this study with $h_{n}=(y_{(n)}-y_{(1)})/(1+\log_{2}n)$.
  5. A5.
    Smoothness. We require the derivative $f^{\prime}(y;\vartheta)$ to be in $L_{2}$. Because $\sigma(t)\geq c>0$, the Gaussian kernel is smooth with bounded derivatives. We differentiate under the integral sign:
    $f^{\prime}(y;\vartheta)\;=\;\int_{0}^{1}\frac{\partial}{\partial y}\phi(y;\mu(t),\sigma^{2}(t))\,g_{T}(t;\psi)\,dt.$
    Recall that $\frac{\partial}{\partial y}\phi(y;\mu,\sigma^{2})=-\frac{y-\mu}{\sigma^{2}}\phi(y)$. The $L_{2}$ norm of this derivative for a single Gaussian component is proportional to $\sigma^{-3/2}$. Since $\sigma(t)\geq c$, there exists a uniform bound $M<\infty$ such that $\|\phi^{\prime}(\cdot|t)\|_{2}\leq M$ for all $t\in[0,1]$. By Minkowski’s inequality for integrals:
    $\|f^{\prime}(\cdot;\vartheta)\|_{2}\;\leq\;\int_{0}^{1}\|\phi^{\prime}(\cdot|t)g_{T}(t;\psi)\|_{2}\,dt\;=\;\int_{0}^{1}g_{T}(t;\psi)\|\phi^{\prime}(\cdot|t)\|_{2}\,dt.$
    Using the uniform bound $M$:
    $\|f^{\prime}(\cdot;\vartheta)\|_{2}\;\leq\;\int_{0}^{1}g_{T}(t;\psi)M\,dt\;=\;M.$
    Thus, $\sup_{\vartheta\in\vartheta}\|f^{\prime}(\cdot;\vartheta)\|_{2}<\infty$, satisfying the smoothness condition.
Appendix C Additional plots
Figure 9: Apical membrance antigen 1 (AMA1): Histograms and fitted Gaussian mixture models for each age group shown in the panels. The means and variances of the two Gaussian components are estimated using data from the age group <6<6 years and are then held fixed for the remaining age groups, while the mixing probabilities are re-estimated.
Figure 10: Merozoite surface protein 1 (MSP1): Histograms and fitted Gaussian mixture models for each age group shown in the panels. The means and variances of the two Gaussian components are estimated using data from the age group <6<6 years and are then held fixed for the remaining age groups, while the mixing probabilities are re-estimated.